OpenStudy (anonymous):

Find the formula, in terms of K, for the area of a regular octagon with apothem K and side of 10.

OpenStudy (anonymous):

A = 1/2*P*a (P is perimeter, a is apothem) A = 1/2(8*10)*K A = 1/2(80)*K A = 40*K A/40 = K

OpenStudy (anonymous):

so its gonna be 40